# Invariant subspaces and PCA in nearly matrix multiplication time

### v3

@@ Abstract
Approximating invariant subspaces of generalized eigenvalue problems (GEPs) is a fundamental computational problem at the core of machine learning and scientific computing. It is, for example, the root of Principal Component Analysis (PCA) for dimensionality reduction, data visualization, and noise filtering, and of Density Functional Theory (DFT), arguably the most popular method to calculate the electronic structure of materials. For a GEP $\boldsymbol{\mathrm{H}}\boldsymbol{\mathrm{C}}=\boldsymbol{\mathrm{S}}\boldsymbol{\mathrm{C}}\boldsymbol{\mathrm{\Lambda}}$, let $\boldsymbol{\mathrm{\Pi}}_{k}$ be the true spectral projector on the invariant subspace that is associated with the $k$ smallest (or largest) eigenvalues. Given $\boldsymbol{\mathrm{H}},$ $\boldsymbol{\mathrm{S}}$, an integer $k$, and accuracy $\epsilon\in(0,1)$, we show that we can compute a matrix $\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}$ such that $\|\boldsymbol{\mathrm{\Pi}}_{k}-\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}\|_{2}\leq\epsilon$, in $O\left(n^{\omega+\eta}\operatorname{polylog}(n,\epsilon^{-1},\kappa(\boldsymbol{\mathrm{S}}),\operatorname{gap}_{k}^{-1})\right)$ bit operations in the floating point model with probability $1-1/n$. Here, $\eta>0$ is arbitrarily small, $\omega\lesssim 2.372$ is the matrix multiplication exponent, $\kappa(\boldsymbol{\mathrm{S}})=\|\boldsymbol{\mathrm{S}}\|_{2}\|\boldsymbol{\mathrm{S}}^{-1}\|_{2}$, and $\operatorname{gap}_{k}$ is the gap between eigenvalues $k$ and $k+1$. To the best of our knowledge, this is the first end-to-end analysis achieving such “forward-error” approximation guarantees with nearly $O(n^{\omega+\eta})$ bit complexity, improving classical $\widetilde{O}(n^{3})$ eigensolvers, even for the regular case $(\boldsymbol{\mathrm{S}}=\boldsymbol{\mathrm{I}})$. Our methods rely on a new $O(n^{\omega+\eta})$ stability analysis for the Cholesky factorization, and a new smoothed analysis for computing spectral gaps, which can be of independent interest. Ultimately, we obtain new matrix multiplication-type bit complexity upper bounds for PCA problems, including classical PCA and (randomized) low-rank approximation.
@@ 1 Introduction
In data science and machine learning, invariant subspaces play a central role in many problems, including Spectral Clustering [NJW01, SM00], Language Models [HysW+22], Image Processing [RAD14, BWS04], Recommendation Systems [DKR02], Principal Components Analysis (PCA) [DK96, Jol02, SSM98], Support Vector Machines [MW05], and many others [FSS20, CEM+15, MG04, BH89]. We particularly focus on PCA applications, which can take the form of a GEP as in Eq. (1) where $\boldsymbol{\mathrm{H}}$ is the sample covariance and $\boldsymbol{\mathrm{S}}$ the identity. In more advanced settings, $\boldsymbol{\mathrm{H}}$ and $\boldsymbol{\mathrm{S}}$ can be defined over a kernel [SSM98, BGAJ11]; See Section 4 and Appendix G for more details. Another closely related application comes from Density Functional Theory [KS65] (DFT), which is not a machine learning problem per se, but it is probably the most commonly used method (it was awarded the Nobel prize in Chemistry in 1998) to compute the electronic and structural properties of materials. In this case, $\boldsymbol{\mathrm{H}}$ is the Hamiltonian and $\boldsymbol{\mathrm{S}}$ the overlap matrix (cf. Appendix F). The spectral projector on the invariant subspace corresponding to the smallest generalized eigenvalues (occupied energies) directly provides the density matrix and the electron density. Obtaining them often presents a challenge from the computational point of view.
@@ 1.1 Problem definition
The complexity of numerical algorithms is often measured in terms of the arithmetic operations executed, commonly referred to as arithmetic complexity. A more realistic notion is the bit complexity, which bounds the number of boolean operations. In the floating point model, it is straightforward to translate the arithmetic to the bit complexity if we have an upper bound on the number of bits. For instance, arithmetic operations on $b$ bits can be typically carried out in $O(b^{2})$ bit operations, or even faster by using more advanced algorithms [SS71, Für07, HVDH21].
@@
In two seminal works [DDH07, DDHK07], it was demonstrated that matrix multiplication and other fundamental problems in Numerical Linear Algebra can be solved in the floating point model with nearly $O(n^{\omega+\eta})$ bit-complexity (up to polylogarithmic factors), where $\eta$ is an arbitrarily small positive number and $\omega$ is the matrix multiplication exponent, to-date bounded to $\omega\lesssim 2.372$ [DWZ23, WXXZ24]. Hereafter, we will use the notation $T_{\mathsf{MM}}(n)=n^{\omega+\eta}$.
@@ 1.2 Existing algorithms
Problem 1.1 can be solved using classic (generalized) eigensolvers and related techniques in $\widetilde{O}(n^{3})$ floating operations, e.g., by reducing the matrix (or pencil) to tridiagonal form with similarity transformations and applying the shifted QR algorithm on the tridiagonal matrix (see [DHT01, NH13, Ips97, BD91, BD98, GVL13, Dem97] and references therein). Significant progresses beyond the $\widetilde{O}(n^{3})$ bit complexity barrier have been made with respect to algorithms for central problems in numerical linear algebra [PC99, DDHK07, DDH07, BGVKS22, LV16, BOE18, DDS23, MMS18]. Still, to the best of our knowledge, the state-of-the-art algorithms providing an end-to-end, provably accurate analysis (in the sense of Eq. (2)) for arbitrary invariant subspaces of definite GEPs or even regular Hermitian eigenproblems, in any finite precision model, remain the classic aforementioned $\widetilde{O}(n^{3})$ algorithms. [MMS18] analyzed the stability of the Lanczos method, and obtained $\widetilde{O}(n^{2})$ bit complexity to compute the top-1 singular vector. The recent pioneering work of [BGVKS22] provided the first $O(T_{\mathsf{MM}}(n))$-type algorithm for the related problem of backward-approximate non-Hermitian matrix diagonalization. It was recently extended to matrix pencils in [DDS23]. However, an end-to-end forward-error analysis of invariant subspace approximations was left as an open problem. Our final results indicate that our main algorithm, which completely avoids diagonalization, is both faster than extending [BGVKS22, DDS23] to solve Problem 1.1, and it requires much less “randomness.” More details about the related literature are provided in Appendix A.7.
@@ 1.3 Contributions and methods
Our main contribution, summarized in the following theorem, is an end-to-end analysis to solve Problem 1.1 in nearly $O(T_{\mathsf{MM}}(n))$ time, thus improving the aforementioned $\widetilde{O}(n^{3})$ type algorithms.
@@
Let $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$ be a Hermitian definite pencil of size $n$ with $\|\boldsymbol{\mathrm{H}}\|,\|\boldsymbol{\mathrm{S}}^{-1}\|\leq 1$, $\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}$ its eigenvalues, $\operatorname{gap}_{k}=\lambda_{k+1}-\lambda_{k}$ and $\kappa(\boldsymbol{\mathrm{S}})=\|\boldsymbol{\mathrm{S}}\|\|\boldsymbol{\mathrm{S}}^{-1}\|$. There exists an algorithm
@@
which takes as inputs $\boldsymbol{\mathrm{H}}$, $\boldsymbol{\mathrm{S}}$, an integer $11\leq k\leq n-$ 1, an error parameter $\epsilon\in(0,1)$ and returns a matrix $\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}$ such that
@@
using the analysis of [BGVKS22] for the Newton iteration. The matrix $\frac{1}{2}(\boldsymbol{\mathrm{I}}+\operatorname{sgn}(\widetilde{\mu}_{k}-\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}}))$ indeed transforms in exact arithmetic all eigenvalues that are smaller than $\widetilde{\mu}_{k}$ to $1$ and the ones larger than $\widetilde{\mu}_{k}$ to zero. As will be proved in Proposition 2.1, in Section 2, this approach is sufficient to provide an accurate spectral projector $\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}$ in floating point. As a consequence, the problem reduces to approximating the aforementioned midpoint and gap. As a baseline, in Appendix B.1 we prove that this can be done in nearly $O(T_{\mathsf{MM}}(n))$ with iterative inversion [DDH07] and diagonalization [BGVKS22] or, similarly, by iteratively calling generalized diagonalization [DDS23]. However, this approach presents two drawbacks: It does not take advantage of the inherent symmetry of the problem, and, at the same time, it performs a full diagonalization when we are only interested in the gap between two specific eigenvalues, which is seemingly redundant. We formally prove this claim by designing a novel approach that achieves better complexity, typically by a factor of $O(\log(n))$ (cf. Section 3.4). Importantly, no explicit diagonalization is necessary.
To minimize the complexity of our algorithm, it is crucial to leverage symmetry. To that end we use the Cholesky factorization of $\boldsymbol{\mathrm{S}}$ in the spirit of the Cholesky-QR algorithm [DHT01]. We highlight that, while other factorizations have been solved in $O(T_{\mathsf{MM}}(n))$ in floating point, an end-to-end analysis for Cholesky remains open. In exact arithmetic, for example, the LU of a Hermitian definite matrix directly provides its Cholesky and [DDH07] showed that the LU factorization of non-symmetric matrices can be obtained in $O(T_{\mathsf{MM}}(n))$. However, when considering arithmetic errors, the relationship between LU and Cholesky does not hold in floating point, as demonstrated by the counter-example of Appendix C.5. Other fast Cholesky algorithms have been proposed for special classes of matrices, e.g., for matrices with well-defined separators [Geo73, LRT79, GT86] and graph Laplacians [KLP+16, KS16]. However, they do not generalize to arbitrary dense matrices. Our analysis is the first to improve the classic $O(n^{3})$ floating point Cholesky algorithms [Kie87, Hig09] for the general case, with provable error bounds. In the following Theorem 1.2 we summarize our new analysis of the Cholesky factorization Algorithm 2 (see Appendix C). We note that the algorithm itself is not new, only its analysis.
@@
This stand-alone result fulfills the definition of “logarithmic-stability,” a notion of numerical stability that is commonly used in the related literature [DDH07, BGVKS22].
Given this new Cholesky analysis, the following transformation of the GEP to a regular Hermitian eigenvalue problem:
@@
The symmetry induced by the Cholesky transformation is crucial to design an efficient algorithm for the spectral gap. As described in Section 3 and analyzed in Appendices D and E, any spectral gap or eigenvalue of a Hermitian definite pencil can be approximated by an iterative algorithm that uses only “counting-queries”, i.e., queries that ask how many eigenvalues are smaller than a given threshold. This can be seen as a generalization of the “PSD-query” algorithm of [LV16] which was used to compute the largest eigenvalue of a matrix. This way we completely avoid diagonalization to compute the gap, thus leading to a lower complexity.
To perform the counting queries efficiently, the transformed matrix $\widetilde{\boldsymbol{\mathrm{H}}}$ must be regularized with small random perturbations, in the spirit of smoothed analysis [ST04], which has recently drawn attention in the context of matrix algorithms [BWZ16, BOE18, BGVKS22, DDS23, MMM24] (see Appendix D for the analysis). These aforementioned works typically require a guarantee on the minimum eigenvalue gap, e.g., [MMM24] uses a Minami-type bound [Min96], while in [BGVKS22, DDS23] the entire pseudospectrum of the perturbed matrix must be shattered with respect to a grid. The latter is even more challenging to achieve than a minimum gap and it requires $\widetilde{O}(n^{2})$ random bits. Our algorithm is significantly less demanding in terms of randomness: All we need is the Wegner estimate [Weg81] for the density-of-states of random Hermitian operators, and only $\widetilde{O}(n)$ random bits in total.
Finally, in Section 4, we apply our main results to prove the first matrix multiplication-type upper bounds for the bit complexity of PCA algorithms. Specifically, for the standard PCA formulation, we show that we can first compute the spectral projector and then use deflation to obtain a basis for the desired low-dimensional embedding in nearly matrix multiplication time. We then apply similar arguments to the seminal Block-Krylov PCA algorithm of [MM15].

### v4

@@ Abstract
Approximating invariant subspaces of generalized eigenvalue problems (GEPs) is a fundamental computational problem at the core of machine learning and scientific computing. It is, for example, the root of Principal Component Analysis (PCA) for dimensionality reduction, data visualization, and noise filtering, and of Density Functional Theory (DFT), arguably the most popular method to calculate the electronic structure of materials. Given Hermitian $\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}}\in\mathbb{C}^{n\times n}$, where $\boldsymbol{\mathrm{S}}$ is positive-definite, let $\boldsymbol{\mathrm{\Pi}}_{k}$ be the true spectral projector on the invariant subspace that is associated with the $k$ smallest (or largest) eigenvalues of the GEP $\boldsymbol{\mathrm{H}}\boldsymbol{\mathrm{C}}=\boldsymbol{\mathrm{S}}\boldsymbol{\mathrm{C}}\boldsymbol{\mathrm{\Lambda}}$, for some $k\in[n]$. We show that we can compute a matrix $\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}$ such that $\|\boldsymbol{\mathrm{\Pi}}_{k}-\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}\|_{2}\leq\epsilon$, in $O\left(n^{\omega+\eta}\operatorname{polylog}(n,\epsilon^{-1},\kappa(\boldsymbol{\mathrm{S}}),\operatorname{gap}_{k}^{-1})\right)$ bit operations in the floating point model, for some $\epsilon\in(0,1)$, with probability $1-1/n$. Here, $\eta>0$ is arbitrarily small, $\omega\lesssim 2.372$ is the matrix multiplication exponent, $\kappa(\boldsymbol{\mathrm{S}})=\|\boldsymbol{\mathrm{S}}\|_{2}\|\boldsymbol{\mathrm{S}}^{-1}\|_{2}$, and $\operatorname{gap}_{k}$ is the gap between eigenvalues $k$ and $k+1$. To achieve such provable “forward-error” guarantees, our methods rely on a new $O(n^{\omega+\eta})$ stability analysis for the Cholesky factorization, and a smoothed analysis for computing spectral gaps, which can be of independent interest. Ultimately, we obtain new matrix multiplication-type bit complexity upper bounds for PCA problems, including classical PCA and (randomized) low-rank approximation.
@@ 1 Introduction
In data science and machine learning, invariant subspaces play a central role in many problems, including Spectral Clustering [111, 124], Language Models [73], Image Processing [118, 11], Recommendation Systems [46], Principal Components Analysis (PCA) [45, 79, 122], Support Vector Machines [99], and many others [50, 35, 96, 12]. We particularly focus on PCA applications, which can take the form of a GEP as in Eq. (1) where $\boldsymbol{\mathrm{H}}$ is the sample covariance and $\boldsymbol{\mathrm{S}}$ the identity. In more advanced settings, $\boldsymbol{\mathrm{H}}$ and $\boldsymbol{\mathrm{S}}$ can be defined over a kernel [122, 20]; See Section 4 and Appendix G for more details. Another closely related application comes from Density Functional Theory [88] (DFT), which is not a machine learning problem per se, but it is probably the most commonly used method (it was awarded the Nobel prize in Chemistry in 1998) to compute the electronic and structural properties of materials. In this case, $\boldsymbol{\mathrm{H}}$ is the Hamiltonian and $\boldsymbol{\mathrm{S}}$ the overlap matrix (cf. Appendix F). The spectral projector on the invariant subspace corresponding to the smallest generalized eigenvalues (occupied energies) directly provides the density matrix and the electron density. Obtaining them often presents a challenge from the computational point of view.
@@ 1.1 Problem definition
The complexity of numerical algorithms is often measured in terms of the arithmetic operations executed, commonly referred to as arithmetic complexity. A more realistic notion is the bit complexity, which bounds the number of boolean operations. In the floating point model, it is straightforward to translate the arithmetic to the bit complexity if we have an upper bound on the number of bits. For instance, arithmetic operations on $b$ bits can be typically carried out in $O(b^{2})$ bit operations, or even faster by using more advanced algorithms [123, 55, 68].
@@
In two seminal works [39, 40], it was demonstrated that matrix multiplication and other fundamental problems in Numerical Linear Algebra can be solved in the floating point model with nearly $O(n^{\omega+\eta})$ bit-complexity (up to polylogarithmic factors), where $\eta$ is an arbitrarily small positive number and $\omega$ is the matrix multiplication exponent, to-date bounded by $\omega\lesssim 2.372$ [47, 141, 6]. Hereafter, we will use the notation $T_{\mathsf{MM}}(n)=n^{\omega+\eta}$.
@@ 1.2 Existing algorithms
Here we give a brief overview of existing algorithms. We refer to Appendix A.7 for more details. GEPs in general can be solved using classic eigensolvers and related techniques in $\widetilde{O}(n^{3})$ floating operations, e.g., by reducing the matrix (or pencil) to tridiagonal form with similarity transformations and applying the shifted QR algorithm on the tridiagonal matrix, or by using a divide-and-conquer method (see [37, 109, 75, 8, 9, 64, 42, 10, 113] and references therein). Significant progresses beyond the $\widetilde{O}(n^{3})$ bit complexity barrier have been made [40, 39, 15, 95, 25, 41, 107, 121]. Regarding the computation of eigenvalues, two notable examples are the $\widetilde{O}(T_{\mathsf{MM}}(n))$ algorithm of [95] for the largest eigenvalye, and the $\widetilde{O}(n^{2})$ algorithm of [107] for the spectral norm.
The first to have addressed the problem of computing invariant subspaces in nearly $O(T_{\mathsf{MM}}(n))$ in floating point is [39] (see also [14, 13]). The authors described an iterative algorithm for the Schur decomposition, and showed that each individual step is numerically stable, and it takes $O(T_{\mathsf{MM}}(n)$ operations. An end-to-end bound on the number of iterations to achieve a backward approximate solution was left open. More recently, the seminal work of [15] extended the analysis to obtain an end-to-end $\widetilde{O}(T_{\mathsf{MM}}(n))$ complexity to approximately diagonalize a matrix, and [41, 121] provided a rigorous analysis for the generalized eigenproblem case. In Corollary 1.7 and Proposition 1.1 of [15], it was also outlined how to translate the backward diagonalization error to a forward error for the eigenvectors. The reported bound, however, has two main limitations: it relies on simplicity of the spectrum, which is a strict assumption, and it requires as input an over-estimate on the eigenvector condition number of the problem, which is unknown, and [15] does not describe how to compute it (see Appendix B.1 for more details). In this work we describe how to overcome these limitations and provide a novel, end-to-end, provably accurate analysis (in the sense of Eq. (2)) for arbitrary invariant subspaces of definite GEPs with $\widetilde{O}(T_{\mathsf{MM}}(n))$ boolean complexity.
@@ 1.3 Contributions and methods
Our main contribution, summarized in the following Theorem 1.1 and Algorithm 1, is an end-to-end analysis to solve Problem 1.1 in nearly $O(T_{\mathsf{MM}}(n))$ time.
@@
Let $(\boldsymbol{\mathrm{H}},\boldsymbol{\mathrm{S}})$ be a Hermitian definite pencil of size $n$ with $\|\boldsymbol{\mathrm{H}}\|,\|\boldsymbol{\mathrm{S}}^{-1}\|\leq 1$, $\lambda_{1}\leq\lambda_{2}\leq\ldots\leq\lambda_{n}$ its eigenvalues, $\operatorname{gap}_{k}=\lambda_{k+1}-\lambda_{k}$ and $\kappa(\boldsymbol{\mathrm{S}})=\|\boldsymbol{\mathrm{S}}\|\|\boldsymbol{\mathrm{S}}^{-1}\|$. Algorithm 1
@@
takes as inputs $\boldsymbol{\mathrm{H}}$, $\boldsymbol{\mathrm{S}}$, an integer $k\in[n-1]$, an error parameter $\epsilon\in(0,1)$ and returns a matrix $\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}$ such that
@@
using the analysis of [15] for the Newton iteration. The matrix $\frac{1}{2}(\boldsymbol{\mathrm{I}}+\operatorname{sgn}(\widetilde{\mu}_{k}-\boldsymbol{\mathrm{S}}^{-1}\boldsymbol{\mathrm{H}}))$ indeed transforms in exact arithmetic all eigenvalues that are smaller than $\widetilde{\mu}_{k}$ to $1$ and the ones larger than $\widetilde{\mu}_{k}$ to zero. As will be proved in Proposition 2.1, in Section 2, this approach is sufficient to provide an accurate spectral projector $\widetilde{\boldsymbol{\mathrm{\Pi}}}_{k}$ in floating point. As a consequence, the problem reduces to approximating the aforementioned midpoint and gap. As a baseline, in Appendix B.1 we prove that this can be done in nearly $O(T_{\mathsf{MM}}(n))$ with iterative inversion [39] and diagonalization [15] or, similarly, by iteratively calling generalized diagonalization [41]. However, this approach presents two drawbacks: It does not take advantage of the inherent symmetry of the problem, and, at the same time, it performs a full diagonalization when we are only interested in the gap between two specific eigenvalues, which is seemingly redundant. We formally prove this claim by designing a novel approach that achieves better complexity, typically by a factor of $O(\log(n))$ (cf. Section 3.4). Importantly, no explicit diagonalization is necessary.
To minimize the complexity of our algorithm, it is crucial to leverage symmetry. To that end we use the Cholesky factorization of $\boldsymbol{\mathrm{S}}$ in the spirit of the Cholesky-QR algorithm [37]. We highlight that, while other factorizations have been solved in $O(T_{\mathsf{MM}}(n))$ in floating point, an end-to-end analysis for Cholesky remains open. In exact arithmetic, for example, the LU of a Hermitian definite matrix directly provides its Cholesky and [39] showed that the LU factorization of non-symmetric matrices can be obtained in $O(T_{\mathsf{MM}}(n))$. However, when considering arithmetic errors, the relationship between LU and Cholesky does not hold in floating point, as demonstrated by the counter-example of Appendix C.5. Other fast Cholesky algorithms have been proposed for special classes of matrices, e.g., for matrices with well-defined separators [58, 94, 61] and graph Laplacians [91, 92]. However, they do not generalize to arbitrary dense matrices. Our analysis is the first to improve the classic $O(n^{3})$ floating point Cholesky algorithms [84, 71] for the general case, with provable error bounds. In the following Theorem 1.2 we summarize our new analysis of the Cholesky factorization Algorithm 2 (see also Appendix C). We note that the algorithm itself is not new, only its analysis.
@@
This stand-alone result fulfills the definition of “logarithmic-stability,” a notion of numerical stability that is commonly used in the related literature [39, 15]. Given this new Cholesky analysis, the following transformation of the GEP to a regular Hermitian eigenvalue problem:
@@
The symmetry induced by the Cholesky transformation is crucial to design an efficient algorithm for the spectral gap. As described in Section 3, and analyzed in Appendices D and E, any spectral gap or eigenvalue of a Hermitian definite pencil can be approximated by an iterative algorithm that uses only “counting-queries”, i.e., queries that ask how many eigenvalues are smaller than a given threshold. This way we completely avoid diagonalization, thus leading to a lower complexity.
To perform the counting queries efficiently, the transformed matrix $\widetilde{\boldsymbol{\mathrm{H}}}$ must be regularized with small random perturbations, in the spirit of smoothed analysis [128], which has recently drawn attention in the context of matrix algorithms [31, 25, 15, 41, 104] (see Appendix D for the analysis). These aforementioned works typically require a guarantee on the minimum eigenvalue gap of the perturbed matrix, e.g., [104] uses a Minami-type bound [105], while in [15, 41] the entire pseudospectrum of the perturbed matrix must be shattered with respect to a grid. The latter is even more challenging to achieve than a minimum gap and it requires $\widetilde{O}(n^{2})$ random bits. Our algorithm is significantly less demanding in terms of randomness: All we need is the Wegner estimate [138, 4] for the density-of-states of random Hermitian operators, and only $\widetilde{O}(n)$ random bits in total.
Finally, in Section 4, we apply our main results to prove the first matrix multiplication-type upper bounds for the bit complexity of PCA algorithms. Specifically, for the standard PCA formulation, we show that we can first compute the spectral projector and then use deflation to obtain a basis for the desired low-dimensional embedding in nearly matrix multiplication time. We then apply similar arguments to the seminal Block-Krylov PCA algorithm of [106].
Algorithm 1 $\mathsf{PROJECTOR}$.
1:
$\mathsf{PROJECTOR}$
8:return $\widetilde{\boldsymbol{\mathrm{\Pi}}}$.
Algorithm 2 $\mathsf{CHOLESKY}(\boldsymbol{\mathrm{M}})$.
1:
$\mathsf{CHOLESKY}$
3:$\boldsymbol{\mathrm{M}}$ is positive-definite, $\textbf{{u}}\leq\epsilon\frac{1}{c_{1}n^{c_{2}}\kappa(\boldsymbol{\mathrm{M}})^{c_{3}\log(n)}}$ for some constants $c_{1},c_{2},c_{3}$ and $\epsilon\in(0,1)$.
5:if $n=1$ :
6:    return $\boldsymbol{\mathrm{L}}=\sqrt{\boldsymbol{\mathrm{M}}_{1,1}}$.
7:else
9:    $\boldsymbol{\mathrm{B}}\boldsymbol{\mathrm{A}}i\leftarrow\mathsf{MM}\left(\boldsymbol{\mathrm{B}},\mathsf{INV}\left(\boldsymbol{\mathrm{A}}\right)\right).$triangleright $\boldsymbol{\mathrm{B}}\boldsymbol{\mathrm{A}}i=\boldsymbol{\mathrm{B}}\left(\boldsymbol{\mathrm{A}}^{-1}+\boldsymbol{\mathrm{E}}_{1}^{\mathsf{INV}}\right)+\boldsymbol{\mathrm{E}}_{2}^{\mathsf{MM}}=\boldsymbol{\mathrm{B}}\boldsymbol{\mathrm{A}}^{-1}+\boldsymbol{\mathrm{E}}_{\boldsymbol{\mathrm{B}}\boldsymbol{\mathrm{A}}i}.$